OpenStudy (megannicole51):

what is the 50th term of the sequence that begins -4, 2, 8, 14?

OpenStudy (reemii):

this sequence seems to be an arithmetic sequence. then can you guess more about it?

OpenStudy (reemii):

basically you need to find the increment and then write the formula for the n-th term. (replace n wiht 50)

OpenStudy (ja1):

so whats the rule? +6?

OpenStudy (megannicole51):

i know theres probably some formula but i dont know what it is

OpenStudy (anonymous):

The sequence begins with -4. There is a common difference of +6 between each number. To find any term in an arithmetic sequence use this. \[a _{n} =a _{0} + (Deltax)(n-1)\]

OpenStudy (anonymous):

\[a _{0} = 4; Deltax = 6; n = 50\]

OpenStudy (reemii):

i think your \(a_0\) should be \(a_1\), or \(n-1\) should be \(n\).

OpenStudy (anonymous):

oops I mean -4

OpenStudy (megannicole51):

(a) 290 (b) 294 (c) 296 (d) 300 here are the answer choices. i'm trying to help my brother with his algebra but i dont remember how to do it and i dont understand the formula you gave

OpenStudy (anonymous):

a

OpenStudy (megannicole51):

thank you:)

OpenStudy (reemii):

to go from one number of the sequence to the next number \(a_k\to a_{k+1}\), in an arithmetic sequence, you just add a certain constant number \(C\). then it's again +C for the next one, +C for the next one , etc. So you add n times the number C if you want to find a number that is at a distance n of the current number. starting number: \(a_1=-4\). number to find: \(a_{50}\). number of steps: 49. therefore, \(a_{50} = a_1 + 49 \times 6 = -4 + 294 = 290\), as DYHW said.

OpenStudy (megannicole51):

thank you that makes a lot of sense:) gotta love formulas lol

OpenStudy (reemii):

there are so many <3

OpenStudy (megannicole51):

why did you multiply 49 by 6?

OpenStudy (megannicole51):

oh nevermind figured it out

OpenStudy (reemii):

ok :D

OpenStudy (megannicole51):

so if i had this question "what is the 20th term of the sequence that begins -3, 6, -12, 24" it would be-3+49*___?

OpenStudy (reemii):

It is not an arithmetic sequence. (you don't add the same value at each step). but there's another "constant operation". if you see what it is you can write , with the same method, the formula for a number at distance 'n' of some number.

OpenStudy (reemii):

do you see what happens at each step? seeing this will unlock everything..

OpenStudy (megannicole51):

well i know its multiplying by -2

OpenStudy (reemii):

that's it, then starting from a value \(x\) and making n steps will get you to the value \(x \times (-2)^n \).

OpenStudy (megannicole51):

so for n do i plug in 20? and im not sure what to plug in for x either

OpenStudy (reemii):

first term (\(a_1\)) of the sequence is a good choice for \(x\), n is the number of steps to get to \(a_{20}\).

OpenStudy (megannicole51):

so x= -3

OpenStudy (reemii):

yes, and the number of steps is 19 (not 20.. ). You'll need a computer to find the value of \(a_{20}\) but you can write the formula for it.

OpenStudy (reemii):

( we wrote everything already) \(a_{20} = -3\times (-2)^{19}\).

OpenStudy (reemii):

explained*

OpenStudy (megannicole51):

awesome thank you:)) uve been a huge help!!

OpenStudy (reemii):

:D